# Linear-time classical approximate optimization of cubic-lattice classical spin glasses

### v4

@@ III Results
This model is described by Eq. (1) with chosen from a uniform distribution over . We use systems with dimensions spins and periodic boundary conditions. To compute the relative energy error from the ground state, we use the value of the ground state energy density for the thermodynamic limit that is numerically estimated in Ref. [32].
We first test the solution quality of the algorithm as a function of fragment size: we apply the bootstrapping algorithm to ten instances of with cubic fragments of size spins with and ; with we find the computation time of exactly contracting a single fragment to be impractically long. The results are shown in Fig. 3; as expected, the error decreases monotonically with , so we use cubic fragments with for the rest of the computations.
For and (ten instances each), the energy error data is shown in Fig. 4, and the time-to-solution (TTS) data is shown in Fig. 5. The minimum error over the tested values of is . The error is non-monotonic in due to finite numerical precision and possibly also finite fragment size. The TTS data closely follows the theoretical expectation of linear scaling.
@@
This model is characterized in Ref. [33]. The model contains multiple base classes of instances, of which we restrict ourselves to the one called , which is shown in Ref. [33] to be computationally the hardest (i.e., taking the longest time to exactly optimize with Markov-Chain Monte-Carlo methods). Ref. [33] also shows that instances in are a few orders of magnitude computationally harder than the cubic-lattice model. We generate instances with the Chook library [34]. For all instances, we enable the option in the Chook library to scramble the ground states with gauge transformations.
@@
As with the model, we use cubic fragments of size 555 and ten problem instances at each value of . The energy error data is shown in Fig. 6, and the TTS data is shown in Fig. 7. The minimum error over the tested values of at is . The error is non-monotonic in due to finite numerical precision and possibly also finite fragment size. The TTS data closely follows the theoretical expectation of linear scaling.
III.3 unweighted Max-Cut on random 3-regular graphs
The seminal work of Barahona in Ref. [4] contains a reduction from the problem of unweighted Max-Cut on random 3-regular graphs to the ground-state problem of Eq. (1) on a simple cubic lattice that is only one cube thick with open boundaries and . The cubic-lattice model preserves the nonlocal correlations of the original problem by having correlated . The reduction has a quadratic overhead () in the number of vertices () of the random graph. We generate random 3-regular graphs with and with the NetworkX library [35], and convert them into cubic-lattice Ising-spin-glass problems via Barahona’s reduction. We then apply the bootstrapping algorithm with a fixed fragment size of 552 spins. We compute energy errors by using the exact ground state energies obtained with the McGroundstate server [36]. We use the results in Ref. [4] to compute the approximation ratio from the energies of the low-energy and ground-state spin configurations.
The energy-error, approximation-ratio, and TTS data are shown in Figs. 8, 9, and 10, respectively. The minimum error over the tested values of is , and the highest approximation ratio for any given instance is at least about 0.72. The error is non-monotonic in due to finite numerical precision and possibly also finite fragment size. The TTS data closely follows the theoretical expectation of linear scaling in , which translates into a quadratic scaling in .
IV Theoretical implications for quantum advantage
The two sources of error in the present heuristic are finite fragment size and finite . The latter can be eliminated by switching from regular tensors to tensors that obey tropical algebra [37], which would allow exact ground state computations within each fragment. (Such a maneuver would block implementation of the algorithm on specialized photonic matrix-multiplication hardware, but is useful for the purely theoretical purpose in this section.) We then observe that for spin glasses with both finite-range couplings and finite-range correlations between the couplings, the average energy error of the modified heuristic would monotonically decay with increasing fragment size. Then, an arbitrarily large but fixed fragment size would yield an arbitrarily small average energy error but still result in a linear time complexity. Therefore, the possibility of quantum advantage in the time complexity of heuristic approximate optimization of such spin glasses is not possible for quantum algorithms with linear or superlinear time complexity, regardless of how rugged the energy landscape of such spin glasses may be.
We note, however, that quantum advantage in the time complexity of heuristic approximate optimization is not the only type of possible quantum advantage, even for heuristic quantum algorithms. It remains a possibility, for example, that heuristic quantum algorithms (such as quantum annealing) may exhibit a quantum advantage in obtaining a diverse range of low-energy solutions within some energy error threshold for classical spin glasses, even with linear or superlinear time complexity.
V Summary and Discussion
We have presented and tested a linear-time, tensor-network heuristic for the approximate optimization of cubic-lattice classical spin glasses. The algorithm consists of approximate computations of single-spin marginals at large inverse temperature via exact contractions of tensor-network fragments; the spins are decimated according to the marginals, and bootstrapping the marginal computations by overlapping the fragments with previously decimated spins yields a global low-energy configuration. We emphasize that the guarantee of linear time complexity comes at the cost of losing any guarantee of solution quality; changing the fragment size, the bootstrapping sequence, the value of , or the numerical precision can yield different solutions, but there is no guarantee on the error of any of them. We empirically found minimum energy errors relative to the ground state of at most a few percent across three different problem classes of the cubic-lattice classical Ising model at the largest sizes that we tested.
The first problem class that we tested was the model used in recent D-Wave experiments. The quantum annealing D-Wave experiments in Ref. [5] were at a fixed size of over 5,000 spins, and they achieved a minimum energy error of about in a time of about 1 . The hybrid quantum-classical D-Wave experiments at a size of over 11,000 spins in Ref. [6] achieved a minimum energy error of about 0.1–1% in about one second of total quantum processing time. In comparison, our classical heuristic consistently yielded a minimum energy error of about at sizes 8,000–125,000 spins in a time – seconds.
The second problem class that we tested was one with an extremely rugged energy landscape: the instance class of the cubic-lattice tile planting model. Ref. [33] showed this problem class to have an exact-optimization time a few orders of magnitude larger than the model for the leading Monte Carlo heuristics at a system size of 888 spins. Here our heuristic consistently yielded a minimum energy error of about at 8,000 spins and about at 125,000 spins; the error monotonically decreased with system size, suggesting extra error at the system boundaries compared to the bulk. The TTS was the same as with the model: – seconds.
The third problem class that we tested was the “two-level spin glass” of Barahona [4], which is a local (cubic-lattice) spin glass that encodes the non-local problem of unweighted Max-Cut on random 3-regular graphs. For this problem class our heuristic consistently yielded a minimum energy error of at system sizes ranging 7,500–270,000 spins, corresponding to random-graph sizes of between 50 and 300 vertices. The approximation ratio was about 0.72–0.88, and the TTS was – seconds. The time-complexity of our deterministic heuristic was verified to be quadratic in the number of vertices. In comparison, the well-known Goemans-Williamson approximation algorithm for Max-Cut [38], which often serves as a benchmark for other Max-Cut algorithms, has a guaranteed approximation ratio of and runs in polynomial time.
While the solution quality of our heuristic on the random-3-regular-graph Max-Cut problem does not approach the typical solution quality of the Goemans-Williamson algorithm, it is not very poor either. This seems surprising given that our heuristic operates via a series of local approximate optimizations, whereas the problem is intrinsically nonlocal. We propose the following explanation: the bootstrapping nature of the algorithm allows the non-local correlations to propagate over the lattice as the spins are successively decimated. It would be interesting in future work to test how well our heuristic works for other local spin-glass problems that encode non-local graph problems. We conjecture that, for fixed fragment sizes, solution quality would decrease with increasing density of the original graph.
On the theoretical side, we explained how the present algorithm implies that quantum advantage in the time complexity of heuristic approximate optimization of classical spin glasses that have decaying correlations is impossible without the candidate quantum algorithm achieving sublinear time complexity, regardless of the ruggedness of the energy landscape of such spin glasses. This disproves the claim in Ref. [12] of a demonstration of quantum advantage in the time complexity of quantum annealing for heuristic approximate optimization of a short-range spin glass with uncorrelated couplings because the quantum algorithm in Ref. [12] exhibits only superlinear time complexity. This also casts substantial doubt on the belief that short-range classical spin glasses with very rugged energy landscapes are good candidates for demonstrating time-complexity quantum advantage in heuristic approximate optimization.
@@ V Summary and Discussion
In industrially-relevant combinatorial optimization problems, what is often desired is a diverse set of low-cost solutions, and some works have therefore addressed the problem of how to sample the low-energy configuration space without bias [39, 40, 41, 42]. We expect that our heuristic will not demonstrate good performance in this regard due to the locality of the single-fragment subroutine and also the enhanced sensitivity to finite numerical precision that arises from the exponential form of the Boltzmann weights that are intrinsic to our heuristic. This is consistent with the fact that in none of the tested problem instances was our heuristic able to obtain the ground state. This expectation is also consistent with the findings in the recent work of Ref. [43]. In that work, a branch-and-bound search strategy was combined with tensor-network-contraction marginal computations to approximately optimize quasi-two-dimensional classical spin lattices, and they found that the number of diverse solutions with less than 1% energy error that was generated with that method was a few orders of magnitude less than the number of such solutions obtained with other methods. For future work, therefore, we propose to hybridize the present method with other methods such that the present method is used to efficiently generate a few low-energy solutions that are in turn used as warm starts for other methods that have better performance in terms of exploring the low-energy configuration space, such as the algorithms in Refs. [40, 42].
@@
We are hopeful that such improvements of the present algorithm may provide a new avenue for gaining insights into the properties of classical spin-glass models studied in condensed matter. For example, questions remain open regarding the nature of the low-temperature phase of the cubic-lattice classical model [47]. Pure Monte Carlo methods have been the state-of-the-art approach for this problem, but the hybrid methods that we propose would operate according to very different principles and may thereby yield new insights. As another example, Ref. [33] points out that the tunability of the tile-planted-solution model can allow for a systematic study of the interplay between disorder and frustration; the algorithms that we propose may provide a complementary route to pure Monte Carlo methods in such a study as well.
We explained that the algorithm may be parallelized through simultaneous use of multiple fragments. This would very substantially reduce the simulation time of our heuristic from what is reported here, which would allow the approximate optimization of much larger systems in a practical amount of time. We also noted that one of the sources of error in our implementation of the algorithm, finite temperature, could be eliminated without changing the time complexity: the ordinary tensors could be replaced by tropical tensors [37]. This would likely reduce the energy errors reported in this work at the same fragment sizes, but would not be compatible with adapting the algorithm to specialized photonic hardware in the manner discussed next.
@@
The computational cost of the present heuristic is dominated by matrix multiplications, for which specialized photonic hardware is under active development. Implementation of a heuristic of the present type with such hardware could yield substantially lower time and energy costs. However, it is unlikely that such hardware will be able to achieve beyond about 8 bits of numerical precision in the foreseeable future [48], whereas the results presented here were all with 64 bits of precision. Our (unshown) preliminary tests of our algorithm with float16 showed very poor results for cubic-lattice tile planting but good results for Barahona’s two-level spin glass. Whether or not the present heuristic can be modified to generally yield sufficiently good warm starts when limited to 8-bit precision is an important open question.
@@
As mentioned, it is possible (and in our opinion very likely) that the heuristic presented in this work will not work well for short-range spin glasses that are reductions or embeddings of dense-graph combinatorial optimization problems. Besides, such optimization problems are more economically represented as dense-graph spin glasses. A tensor-network algorithm that can efficiently optimize dense-graph rugged-energy-landscape spin glasses is therefore desirable. Ref. [49] demonstrates that all-to-all coupled Ising spin glasses can sometimes be optimized by iterating over optimizations of randomly selected subsets of all-to-all coupled spins. In future work we plan to investigate such an approach with the tensor-network machinery that we have demonstrated here, such that the subsets of all-to-all coupled spins take the place of the local fragments in the present algorithm. Our (unshown) preliminary tests on small but very rugged instances of the all-to-all coupled Wishart planted ensemble [50] show good results for approximate optimization without decomposition into subsets. As with the heuristic that we demonstrated in the present work, such a tensor-network approach would be straightforward to generalize to beyond two-body interactions and beyond binary variables.

### v6

@@ III Experiments
This model is described by Eq. (1) with chosen from a uniform distribution over . We use systems with dimensions spins and periodic boundary conditions. The relative energy error from the ground state is given by
where is the Hamiltonian energy of the computed configuration and is the ground state energy. We compute this by using the value of the ground state energy density for the thermodynamic limit that is numerically estimated in Ref. [65].
We first test the solution quality of the algorithm as a function of fragment size: we apply the bootstrapping algorithm to ten instances of with cubic fragments of size spins with and ; with we find the computation time of exactly contracting a single fragment to be impractically long. The results are shown in Fig. 3; as expected from the results in Ref. [66], the error decreases monotonically with , so we use cubic fragments with for the rest of the computations.
For and (ten instances each), the energy error data is shown in Fig. 4. The minimum error over the tested values of is , and the error fluctuation is less than 1%. The error is non-monotonic in due to finite numerical precision and possibly also finite fragment size. The TTS data in Fig. 10 of Appendix B closely follows the theoretical expectation of linear scaling.
@@
This model is characterized in Ref. [31]. The model contains multiple base classes of instances; the three known as , , and can be generated by the Python library called Chook [67]. Following Ref. [31] we generate two families of instances: gallus_26 (a mixture of and ), and gallus_46 (a mixture of and ). Both families are parameterized by , which denotes the fraction of the full-system Hamiltonian that belongs to . For all instances we enable the option in the Chook library to scramble the ground states with gauge transformations. The error is the defined the same as in Eq. (3), except now is given exactly by the Chook library for each generated instance.
Between and , the hardness of exact optimization with MCMC of gallus_26 and gallus_46 is shown in Ref. [31] to increase (roughly) monotonically with increasing . When the MCMC exact-optimization hardness of gallus_46 is equivalent to that of the model, and the MCMC exact-optimization hardness of gallus_26 is slightly higher. When , gallus_26 and gallus_46 are equivalent and have an MCMC exact-optimization hardness a few orders of magnitude greater than the model.
@@
We first test the performance of our heuristic on instances from the class (i.e., ) with different system sizes. As with the model, we use cubic fragments of size 555 and ten problem instances at each value of . The energy error data is shown in Fig. 5 and Fig. 6. The error is non-monotonic in due to finite numerical precision and possibly also finite fragment size, but monotonically decreases with increasing system size. Fig. 11 in Appendix B confirms that the TTS data closely follows the theoretical expectation of linear scaling.
From Fig. 6 we estimate an upperbound on the asymptotic error for the linear-time heuristic of 4% for the class, and we independently optimize the same instances from with SA and PT to 4% error. Plotting the data for SA and PT on semilog plots vs. and several smaller powers of (not shown) does not reveal a straight line, so we conclude that the SA and PT scaling is most likely polynomial. Fitting to finds for SA (Fig. 7) and for PT (Fig. 8), compared to 1.0081 for the tensor-network heuristic (Fig. 11). We note that this is even though the data that is used to compute the scalings for SA and PT is from runs with instance-specific tuning of the hyperparameters (see Appendix A). Thus we conclude a slight scaling advantage for the subsystem-based tensor-network heuristic over full-system SA and PT.
We next test our heuristic on gallus_26 and gallus_46 with and . The results are displayed in Fig. 9. We find an (empirical) upper bound on the error of about 7.5%, which is almost a factor of two better than the upper bound of 11.8% [60] known for non-heuristic approximate optimization algorithms for Eq. (1) that operate in polynomial time. We note, however, that this upper bound for non-heuristic approximate optimization takes into account the worst possible cases over all possible geometries for Eq. (1). Comparing to it the upper bound of our heuristic that is specialized to the cubic lattice is therefore not a completely fair comparison, but it is still of interest since it shows that our heuristic at least beats that more general upper bound of non-heuristic approximate algorithms. Interestingly, though Ref. [31] confirms increasing landscape ruggedness (i.e., MCMC hardness for exact optimization) as increases between 0.8 and 1, our linear-time heuristic actually shows a monotonically decreasing error.
IV Summary and Outlook
The question of where quantum speedups lie is an active area of research. It is thought that optimization of classical spin glasses with rugged energy landscapes may be a good place to look for such speedups due to the potential of quantum tunneling. Establishing a quantum speedup for a specific problem requires a comparison with the best known scaling behavior of classical heuristics on that same problem. Markov-Chain-Monte-Carlo (MCMC) algorithms, such as simulated annealing (SA) and parallel tempering (PT), have for a long time been considered the gold standard of classical heuristics for optimization of classical spin glasses with rugged energy landscapes. In this work we showed that a certain non-MCMC classical heuristic can sometimes have superior scaling to SA and PT for approximate optimization of rugged-energy-landscape classical Ising spin glasses on the cubic lattice. This suggests that classical heuristics aside from SA and PT should be taken into account when attempting to establish a quantum speedup on this type of problem, which is relevant to the state-of-the-art of D-Wave’s experimental capability for approximate optimization of classical spin glasses.
The non-MCMC classical heuristic that we introduced is actually one version of a meta-heuristic: instead of optimizing the full problem directly, it concatenates (via a bootstrapping procedure) the results of subsystem optimizations, but the subsystem optimizations can be done, in principle, with any heuristic. We used a tensor-network heuristic for the subsystem optimizations. The linear scaling of the meta-heuristic would not change by using a different heuristic for the subsystem optimizations, but there could be a big difference in the wall time. In principle, the tensor-network subsystem heuristic could outperform (both in terms of wall time and optimization error) other classical heuristics for the subsystem optimizations if the energy landscape of the subsystems was sufficiently rugged because the tensor-network subsystem heuristic does not perform a search over the energy landscape. Thus, the present classical version of the meta-heuristic may be a way to partially overcome the computational hardness of extremely rugged energy landscapes. On the other hand, quantum heuristics that exploit tunneling may also outperform search-based classical heuristics for optimization of subsystems with sufficiently rugged energy landscapes. All-in-all, certain versions of the present meta-heuristic have the potential to alleviate the computational hardness of rugged energy landscapes through two possible ways: 1) replacing full-system optimization with multiple subsystem optimizations, and 2) performing the subsystem optimizations with a classical heuristic that does not perform a search over the energy landscape or with a quantum heuristic that exploits tunneling.
In the present work we only showed a slight scaling advantage of the meta-heuristic over full-system SA and full-system PT. Demonstrating a stronger scaling advantage over a wider variety of competing heuristics remains an open task. If the present meta-heuristic could be modified to yield lower error, this may result in a stronger scaling advantage. This might be achieved by the following form of post-processing on the output of the present meta-heuristic: undecimate the spins in a randomly selected fragment, approximately optimize that fragment again, accept the new configuration only if it lowers the global energy, and repeat on a new random fragment. It would be straightforward to perform this post-processing in a way that preserves the linear scaling of the overall heuristic, but the scaling of the error in such a case would need to be empirically checked. Also, we note that one of the sources of error in our implementation of the meta-heuristic, finite temperature, could be eliminated without changing the time complexity: the ordinary tensors could be replaced by tropical tensors [68]. This would likely reduce the energy errors reported in this work at the same fragment sizes, however it would not be compatible with adapting the meta-heuristic to specialized photonic hardware in the manner discussed below.
@@ IV Summary and Outlook
In industrially-relevant combinatorial optimization problems, what is often desired is a diverse set of low-cost solutions, and some works have therefore addressed the problem of how to sample the low-energy configuration space without bias [69, 70, 71, 72]. We expect that our heuristic will not demonstrate good performance in this regard due to the locality of the single-fragment subroutine and also the enhanced sensitivity to finite numerical precision that arises from the exponential form of the Boltzmann weights that are intrinsic to our heuristic. This is consistent with the fact that in none of the tested problem instances was our heuristic able to obtain the ground state. This expectation is also consistent with the findings in the recent work of Ref. [73]. In that work, a branch-and-bound search strategy was combined with tensor-network-contraction marginal computations to approximately optimize quasi-two-dimensional classical spin lattices, and they found that the number of diverse solutions with less than 1% energy error that was generated with that method was a few orders of magnitude less than the number of such solutions obtained with other methods. For future work, therefore, we propose to hybridize the present method with other methods such that the present method is used to efficiently generate a few low-energy solutions that are in turn used as warm starts for other methods that have better performance in terms of exploring the low-energy configuration space, such as the algorithms in Refs. [70, 72]. It may be that such a hybridization produces a classical heuristic with an average-case time complexity that strongly challenges that of quantum heuristics for obtaining ground and low-energy states of very rugged energy landscape spin glasses.
@@
We speculate that such improvements of the present algorithm may provide a new avenue for gaining insights into the properties of classical spin-glass models that are of interest in condensed matter. For example, questions remain open regarding the nature of the low-temperature phase of the cubic-lattice classical model [77]. Traditional Monte Carlo methods have been the state-of-the-art approach for this problem, but the hybrid methods that we propose would operate according to very different principles and may thereby yield new insights. As another example, Ref. [31] points out that the tunability of the tile-planted-solution model can allow for a systematic study of the interplay between disorder and frustration; the algorithms that we propose may provide a complementary route to traditional Monte Carlo methods in such a study as well.
@@
The wall time of the present form of our meta-heuristic (i.e., using tensor-network contraction for the subsystem optimizations) does not compare favorably with full-system SA. However, we explained that the meta-heuristic may be parallelized through simultaneous use of multiple fragments. This would very substantially reduce the simulation time of our heuristic from what is reported here. Also, GPU acceleration has been shown to decrease the walltime of tensor-network contraction by over one order of magnitude [78]. Further, the computational cost of tensor-network contraction is dominated by matrix multiplications, for which specialized photonic hardware is under active development. Implementation of tensor-network contraction with such hardware could in principle yield substantially lower time and energy costs. However, it is unlikely that such hardware will be able to achieve beyond about 8 bits of numerical precision in the foreseeable future [79], whereas the results presented here were all with 64 bits of precision. Our (unshown) preliminary tests of our algorithm with float16 showed very poor results for the model and cubic-lattice tile planting model but modest results for Barahona’s two-level spin glass [2], which is a cubic-lattice reduction of Max-Cut on random three-regular graphs. Whether or not the present heuristic can be modified to generally yield sufficiently good warm starts when limited to 8-bit precision is an important open question.
@@
It is possible that the heuristic presented in this work will not work well for short-range spin glasses that are reductions or embeddings of dense-graph combinatorial optimization problems. Besides, such optimization problems are more economically represented as dense-graph spin glasses. A tensor-network algorithm that can efficiently optimize dense-graph rugged-energy-landscape spin glasses is therefore desirable. Ref. [80] demonstrates that all-to-all coupled Ising spin glasses can sometimes be optimized by iterating over optimizations of randomly selected subsets of all-to-all coupled spins. In ongoing work using the Wishart planted ensemble [81], which is an all-to-all coupled classical Ising spin glass model with tunable ruggedness, we are investigating such an approach where the tensor-network machinery that we have demonstrated here is used to approximately optimize the random subsystems. As with the tensor-network-based approach that we demonstrated in the present work, such a tensor-network approach would be straightforward to generalize to beyond two-body interactions and beyond binary variables.